A ball is thrown forward horizontally from the top of a cliff with a velocity 10ms-1. The






height of the cliff above the ground is 45m. Calculate (i) the time to reach the ground (ii)






the distance from the cliff of the ball on hitting the ground (iii) the velocity with which






the ball hits the ground. (iv) the direction of the ball to the horizontal just before it hits






the ground

1
Expert's answer
2022-02-16T08:35:34-0500
"t=\sqrt{2\frac{45}{9.8}}=3.0\ s\\d=(10)(3)=30\ m\\
\theta=\arctan{\frac{9.8\cdot3}{10}}=71\degree\text{ with horizontal}"